# Beyond Grounding: Extracting Fine-Grained Event Hierarchies Across Modalities

### 14 Dec 2023 v2

@@ Abstract
In this paper, we propose the task of extracting event hierarchies from multimodal (video and text) data to capture how the same event manifests itself in different modalities at different semantic levels. This reveals the structure of events and is critical to understanding them. To support research on this task, we introduce the Multimodal Hierarchical Events (MultiHiEve) dataset. Unlike prior video-language datasets, MultiHiEve is composed of news video-article pairs, which makes it rich in event hierarchies. We densely annotate a part of the dataset to construct the test benchmark. We show the limitations of state-of-the-art unimodal and multimodal baselines on this task. Further, we address these limitations via a new weakly supervised model, leveraging only unannotated video-article pairs from MultiHiEve. We perform a thorough evaluation of our proposed method which demonstrates improved performance on this task and highlight opportunities for future research.
@@ 1 Introduction
Multimodal event hierarchies can aid many applications, such as summarization (Daumé III and Marcu 2006), story completion from multiple sources, event analysis/comparison (e.g. a “protest” event with “property destruction” subevent is unruly, otherwise it’s peaceful), event prediction likelihood (Chaturvedi, Peng, and Roth 2017), knowledge-based information extraction (Wen et al. 2021a), and multimodal knowledge graph construction (Li et al. 2020a).
@@
To benchmark performance on this task, we construct several baselines using state-of-the-art (SOTA) architectural components. A unimodal text-only baseline leverages ASR (automated speech recognition) and employs a SOTA NLP model (Wang et al. 2021) to find hierarchical events between a text article and its video’s ASR. We also build a multimodal baseline by detecting hierarchical events in text using (Wang et al. 2021) and grounding the textual subevents to video using CLIP (Radford et al. 2021). A key limitation of these baselines is that they require visual subevents to be mentioned in textual form in either the ASR or the article. To address this limitation, we propose Multimodal Analysis of Structured Hierarchical Events Relations (MASHER), a weakly supervised model which learns to directly predict hierarchical events between text and video. By doing so, MASHER can also discover visual-only subevents (subevents not mentioned in text).
@@ 2 Related Work
Hierarchical Event Relations in Text. Detecting hierarchical event relations (or sub-event relations) is a long-standing problem in the text domain (O’Gorman, Wright-Bettner, and Palmer 2016; Glavaš et al. 2014). Early works mainly rely on heuristic phrasal patterns. For example, Badgett and Huang (2016) found some characteristic phrases (e.g. “media reports” in new articles) always contain sub-events with hierarchical relations. To further enrich hierarchical event relation instances, recent works (Yao et al. 2020a) rely on generative language model to generate subevent knowledge among different commonsense knowledge (Bosselut et al. 2019; Sap et al. 2019), then incorporate knowledge into event ontology. In this paper, we focus on the hierarchical event relations between multi-modal events.
@@
Relation Understanding in the Vision Domain. Prior work (Krishna et al. 2017b; Xu et al. 2017; Ji et al. 2020) propose scene graph methods which parse images/videos into a graph. However, the relationships studied in scene graphs are not between two events. To the best of our knowledge, the only pioneering work that has discussed the event-event relationship in video domain is VidSitu (Sadhu et al. 2021). Unfortunately, to simplify the research problem on this topic, they have made several assumptions: 1) All events are manually cutted into fixed interval (2-second). 2) All event types are “visual” only and from a fixed event ontology. On the other hand, we consider variable-length video events and focus on open-vocabulary event types (which include non-visual event types and other visual events like funeral, detain, rally etc). Besides, while they have annotated each video event with a text label, their event relations still are between events in a video. In contrast, our multimodal relations are between events in an article and a video (see Figure 2).
@@
Multimodal Event Understanding. Since single-modality event extraction is well studied (Nguyen, Cho, and Grishman 2016; Sha et al. 2018; Liu et al. 2019, 2020; Yatskar, Zettlemoyer, and Farhadi 2016b; Li et al. 2017; Mallya and Lazebnik 2017; Pratt et al. 2020a, b; Yatskar, Zettlemoyer, and Farhadi 2016a), understanding events from multiple modalities (Li et al. 2020b; Chen et al. 2021; Reddy et al. 2021; Li et al. 2022; Zhang et al. 2017; Tong et al. 2020; Wen et al. 2021b; Park et al. 2020; Reddy et al. 2022) has attracted extensive research interests because different modalities usually provide the complementary information for comprehensively understanding the real-world complex events. Two important benchmarks (Li et al. 2020b; Chen et al. 2021) have been established for image + text and video + text settings. Li et al. (2020b) first introduced the task of jointly extracting events and labeling argument roles from both text articles and images.  Chen et al. (2021) further defined the task of joint multimedia event extraction from video and text to exploit the rich dynamics from videos. However, both the works focus on event detection in comparison to the event relations task explored in this work.
@@ 3 Task
Text Event Definition. The definition of an “event” has been defined quite thoroughly in different NLP works on information extraction (Huang et al. 2004; Reimers, Dehghani, and Gurevych 2016; GLAVAŠ and ŠNAJDER 2015). As such, we closely follow ACE Corpus’s (Huang et al. 2004) definition of an event: ‘a change of state or the changed state of a person, thing or entity.(Huang et al. 2004) also provides a more detailed criteria addressing linguistic nuances not required in our setting. Secondly, (Huang et al. 2004) is restricted to a fixed ontology of events, whereas we annotate events in an open world. As a result, we came up with a modified event definition and annotation criteria (detailed in Section A.1).
@@
Dataset Domain Text Type Words/min. [^1] For datasets with duplicate descriptions per video clip ((Chen and Dolan 2011; Xu et al. 2016)), words/min. is averaged by #descriptions .
@@ 6.1 Baselines
Text-only Baseline. (Text Base.) To evaluate the contribution of visual data in generating multimodal event hierarchy, we compare our model against a text-only baseline. We construct this baseline by using ASR provided with video as a proxy for video events (c.f. Section D.2). Specifically, the proxy for video event $v_{j}$ is the ASR found within the timestamps of $v_{j}$, denoted $X_{j}$. We extract events from $X_{j}$, $\{e^{\prime}_{j_{z}}\}_{z=1}^{w}$ following Section 5.1. Next, we use the NLP model (Wang et al. 2021) to predict the relationship type, $r_{ij_{z}}$ between the text event from the article, $e_{i}$, and proxy video events from ASR, $e^{\prime}_{j_{z}}$. If any $r_{ij_{z}}\in\{\textrm{``Hierarchical'', ``Identical''}\}\forall z$, we propagate the relation $r_{ij_{z}}$ from $e_{i}$ and $e^{\prime}_{j_{z}}$ to $e_{i}$ and $v_{j}$ because $e^{\prime}_{j_{z}}$ is a proxy for $v_{j}$.
Multimodal Baseline. (MM Base.) We discussed a method to predict multimodal relations in Section 5.1, which used NLP and vision methods to produce pseudo labels. This is currently the best performance that NLP and vision can combine to give without a trained model. As such, we consider this pipeline as our multimodal baseline.
@@ 6.2 Results
  • MASHER achieves 4x recall over MM Base for hierarchical relations. It is because MM Base. relies on finding the subevent in text first before retrieving the matching video subevent (Section 5.1). This causes it to miss visual-only subevents while MASHER can discover those as it directly predicts multimodal relations. This fact is evident in Figure 4 – MM Base can only discover “took” (to streets) video subevent as it is mentioned in text as well. MASHER, in comparison, can also detect “water cannon” visual-only subevent. We further validate this hypothesis by calculating recall on visual-only subevents. MASHER scores 15.92% while MM Base. scores 2.14%. This also explains MM Base’s better precision, since it only predicts a few relations.
@@ C.4 Implementation Details
Appendix D Experimental Details and Setup
@@ D.2 Text Only Baseline Details
D.3 LLM Based Baseline: BLIP-2
We consider a large language based multimodal baseline: BLIP-2 (Li et al. 2023). As BLIP-2 is image based, not video based, we input the middle from the video event as a proxy for the video event. Alogn with this we input text event and prompt it with the definition of the task and the three types of relations. We report the results in Table 10. BLIP-2 fails to predict any relation as Hierarchical. This is not surprising as this is a completely new task. However, BLIP-2 does perform better than Text Base. for Identical relations, demonstrating the value of visual data to this task. Indeed, it’s telling that it performs worse than MM Base. Perhaps, it indicates the superiority of CLIP in multimodal matching tasks.

### 20 Dec 2023 v3

@@ Abstract
In this paper, we propose the task of extracting event hierarchies from multimodal (video and text) data to capture how the same event manifests itself in different modalities at different semantic levels. This reveals the structure of events and is critical to understanding them. To support research on this task, we introduce the Multimodal Hierarchical Events (MultiHiEve) dataset [^1] https://github.com/hayyubi/multihieve. Unlike prior video-language datasets, MultiHiEve is composed of news video-article pairs, which makes it rich in event hierarchies. We densely annotate a part of the dataset to construct the test benchmark. We show the limitations of state-of-the-art unimodal and multimodal baselines on this task. Further, we address these limitations via a new weakly supervised model, leveraging only unannotated video-article pairs from MultiHiEve. We perform a thorough evaluation of our proposed method which demonstrates improved performance on this task and highlight opportunities for future research.
@@ 1 Introduction
Multimodal event hierarchies can aid many applications, such as summarization (Daumé III and Marcu 2006), story completion from multiple sources, event analysis/comparison (e.g. a “protest” event with “property destruction” subevent is unruly, otherwise it’s peaceful), event prediction likelihood (Chaturvedi, Peng, and Roth 2017), knowledge-based information extraction (Wen et al. 2021), and multimodal knowledge graph construction (Li et al. 2020a).
@@
To benchmark performance on this task, we construct several baselines using state-of-the-art (SOTA) architectural components. A unimodal text-only baseline leverages ASR (automated speech recognition) and employs a SOTA NLP model (Wang et al. 2021) to find hierarchical events between a text article and its video’s ASR. We also build a multimodal baseline by detecting hierarchical events in text using Wang et al. (2021) and grounding the textual subevents to video using CLIP (Radford et al. 2021). A key limitation of these baselines is that they require visual subevents to be mentioned in textual form in either the ASR or the article. To address this, we propose Multimodal Analysis of Structured Hierarchical Events Relations (MASHER), a weakly supervised model which learns to directly predict hierarchical events between text and video. By doing so, MASHER can also discover visual-only subevents (subevents not mentioned in text).
@@ 2 Related Work
Hierarchical Event Relations in Text. Detecting hierarchical event relations (or sub-event relations) is a long-standing problem in the text domain (O’Gorman, Wright-Bettner, and Palmer 2016; Glavaš et al. 2014). Early works mainly rely on heuristic phrasal patterns. For example, Badgett and Huang (2016) found some characteristic phrases (e.g. “media reports” in new articles) always contain sub-events with hierarchical relations. To further enrich hierarchical event relation instances, recent works (Yao et al. 2020a) rely on generative language model to generate subevent knowledge among different commonsense knowledge (Bosselut et al. 2019; Sap et al. 2019), then incorporate knowledge into event ontology.
@@
Relation Understanding in the Vision Domain. Prior work (Krishna et al. 2017b; Xu et al. 2017; Ji et al. 2020) propose scene graph methods that parse images/videos into a graph. However, the relationships studied in scene graphs are not between two events. To the best of our knowledge, the only pioneering work that has discussed the event-event relationship in video domain is VidSitu (Sadhu et al. 2021). Unfortunately, to simplify the research problem on this topic, they have made several assumptions: 1) All events are manually cutted into fixed interval (2-second). 2) All event types are “visual” only and from a fixed event ontology. On the other hand, we consider variable-length video events and focus on open-vocabulary event types (which include non-visual event types and other visual events like funeral, detain, rally etc). Besides, while they have annotated each video event with a text label, their event relations still are between events in a video. In contrast, our multimodal relations are between events in an article and a video (see Figure 2).
@@
Multimodal Event Understanding. Since single-modality event tasks are well studied (Nguyen, Cho, and Grishman 2016; Sha et al. 2018; Liu et al. 2019, 2020; Li et al. 2017; Mallya and Lazebnik 2017; Pratt et al. 2020; Yatskar, Zettlemoyer, and Farhadi 2016; Lu et al. 2023), jointly understanding events from multiple modalities (Li et al. 2020b; Chen et al. 2021; Li et al. 2022; Zhang et al. 2017; Tong et al. 2020; Wen et al. 2021; Park et al. 2020; Reddy et al. 2022; Du et al. 2022) has attracted extensive research interests because different modalities usually provide the complementary information for comprehensively understanding the real-world complex events. Two important benchmarks (Li et al. 2020b; Chen et al. 2021) have been established for image + text and video + text settings. Li et al. (2020b) first introduced the task of jointly extracting events and labeling argument roles from both text articles and images.  Chen et al. (2021) further defined the task of joint multimedia event extraction from video and text to exploit the rich dynamics from videos. However, both the works focus on event detection in comparison to the event relations task explored in this work.
@@ 3 Task
Text Event Definition. The definition of an “event” has been defined quite thoroughly in different NLP works on information extraction (Huang et al. 2004; Reimers, Dehghani, and Gurevych 2016; GLAVAŠ and ŠNAJDER 2015). As such, we closely follow ACE Corpus’s (Huang et al. 2004) definition of an event: ‘a change of state or the changed state of a person, thing or entity.’ We came up with a slightly modified event definition and annotation criteria (detailed in Section A.1) as ACE 1) addresses several linguistic nuances not required in our setting; 2) is restricted to a fixed ontology of events, whereas we annotate open-domain events.
@@
Dataset Domain Text Type Words/min. [^2] For datasets with duplicate descriptions per video clip (MSVD, MSR-VTT), words/min. is averaged by #descriptions .
@@ 6.1 Baselines
[…]
Multimodal Baseline. (MM Base.) We discussed a method to predict multimodal relations in Section 5.1, which used NLP and vision methods to produce pseudo labels. This is currently the best performance that NLP and vision can separately combine to give without a trained model. As such, we consider this pipeline as our multimodal baseline.
Video-LLaMA (Zhang, Li, and Bing 2023). It is a video based large language model which has demonstrated strong zero-shot results on multiple video language tasks. Consequently, we consider it as one of our baseline.
@@ 6.2 Results
  • MASHER achieves 4x recall over MM Base for hierarchical relations. It is because MM Base. relies on finding the subevent in text first before retrieving the matching video subevent (Section 5.1). This causes it to miss visual-only subevents while MASHER can discover those as it directly predicts multimodal relations. This fact is evident in Figure 4 – MM Base can only discover “took” (to streets) video subevent as it is mentioned in text as well. While MASHER can also detect “water cannon” visual-only subevent. We further validate this hypothesis by measuring recall on visual-only subevents. MASHER scores 15.92% while MM Base. scores 2.14%. This also explains MM Base’s better precision, since it only predicts a few relations.
@@
  • Video-LLaMA, being a strong multimodal baseline, outperforms both Prior and Text Base. However, it’s worse than MM Base, indicating specialized models (MM Base.) perform better than generic vision-language model (Video-LLaMA) on this task.
@@ C.4 Implementation Details
Appendix D Evaluation Metric and Baselines
@@ D.2 Text Only Baseline Details
D.3 Image-based LLM Baseline: BLIP-2
We consider a large language image-based multimodal baseline: BLIP-2 (Li et al. 2023). This was done to study how image-based LLM can generalize to the proposed video task. As BLIP-2 is image based, not video based, we separately input 5 uniformly sampled frames from the video event and take the majority prediction as output. Along with the frame we input text event and prompt it with the definition of the task and the three types of relations. We report the results in Table 10. BLIP-2 fails to predict any relation as Hierarchical. This is not surprising as this is a completely new task. However, BLIP-2 does perform better than Text Base. for Identical relations, demonstrating the value of visual data to this task. Indeed, it’s telling that it performs worse than MM Base. Perhaps, it indicates the superiority of CLIP in multimodal matching tasks.